### v1

# Anisotropy of the anomalous Hall effect in the altermagnet candidate MnSi films

Abstract
Altermagnets are compensated magnets belonging to spin symmetry groups that allow alternating spin polarizations both in the coordinate space of the crystal and in the momentum space of the electronic structure. In these materials the anisotropic local crystal environment of the different sublattices lowers the symmetry of the system so that the opposite-spin sublattices are connected only by rotations, which results in an unconventional spin-polarized band structure in the momentum space. This low symmetry of the crystal structure is expected to be reflected in the anisotropy of the anomalous Hall effect. In this work, we study the anisotropy of the anomalous Hall effect in epitaxial thin films of MnSi, an altermagnetic candidate material. We first demonstrate a change in the relative Néel vector orientation when rotating the external field orientation through systematic changes in both the anomalous Hall effect and the anisotropic longitudinal magnetoresistance. We then show that the anomalous Hall effect in this material is anisotropic with the Néel vector orientation relative to the crystal structure and that this anisotropy requires high crystal quality and unlikely correlates with the magnetocrystalline anisotropy. Our results provide further systematic support to the case for considering epitaxial thin films of MnSi as an altermagnetic candidate material.
I Introduction
The physical properties of magnetic materials are governed by the intrinsic symmetries of their crystal and spin structures, which are represented by symmetry groups. A good example in the realm of magnetotransport properties is the anomalous Hall effect (AHE) [1], where a longitudinal electric field generates a transversal Hall current that is perpendicular to the Hall vector (, where are the off-diagonal terms of the conductivity matrix) i.e. . As is an axial vector and odd upon time-reversal (), non-vanishing AHE is only observed in systems where the symmetries allow for such a vector. In ferromagnetic materials the net magnetization is also a -odd axial vector and thus is allowed [1] while in magnetic materials with compensated order, is allowed upon sufficient lowering of the crystal symmetries [2, 3, 4, 5, 6, 7]. In antiferromagnetic materials [8, 9] where the opposite-spin sublattices are connected by translation or inversion, is forbidden by symmetry, while in altermagnetic materials [10] this connecting symmetry is rotation, which allows for [11, 12, 13, 14, 15, 16, 17]. This lowered symmetry in altermagnets stems from the two sublattices being inequivalent in terms of the local crystal environment. Altermagnets are characterised by alternating spin-splitting of the energy bands [18, 19, 20, 21, 22] and they show various core spintronic properties such as the AHE presented above, the Nernst effect (thermal counterpart of the AHE) [23], spin-polarized current generation [24, 25, 26], as well as giant and tunnel magnetoresistances [27, 28, 29] in heterostructures.
In polycrystalline soft ferromagnets, is allowed by symmetry for any magnetization orientation. Therefore, the AHE is proportional to the component of the magnetization that is parallel to , i.e. to cos, where is the saturation magnetization and is the angle between the external field and the film normal, assuming that the field is strong enough to saturate the magnetization. Exceptions to this behavior can arise in crystalline ferromagnets [30, 31] or ferromagnets with strong magnetocrystalline anisotropy, which causes a lag between the external field and the magnetization. In altermagnets, specific Néel vector () orientations can either allow or exclude the Hall vector by symmetry [7, 11, 13] so that anisotropic AHE beyond cos can arise.
Anisotropic AHE beyond cos in compensated magnets constitutes a hallmark of altermagnetism, as has been recently demonstrated in RuO [11] (intrinsic d-wave altermagnet, i.e. the electronic structure displays four lobes with alternating spin-polarization) and MnTe showing spontaneous AHE [13] (intrinsic g-wave, which acquires a d-wave character in the presence of spin-orbit coupling). Both RuO and MnTe owe their altermagnetism to the presence of non-magnetic atoms that lower the symmetry of the system such that the opposite-spin sublattices are not connected by either inversion or translation [10]. The saturation of in RuO and MnTe requires high fields of 68 [11, 12] and 4 T [13], respectively, so for further studies on the anisotropy of AHE in altermagnets a new material candidate where -saturation is realized at smaller fields would be ideal.
Recently, epitaxial thin films of MnSi have emerged as such candidate [15] as the collinear compensated phase of MnSi with a hexagonal unit cell and four magnetic Mn atoms on the Mn2 sites in a checkerboard arrangement of the sublattices shown in Figure 1(a) fulfills the symmetry criteria for altermagnetism [15]. Note, that in the bulk phase of MnSi the hexagonal unit cell is not preserved during the magnetic ordering and the opposite-spin sublattices become connected by translation making bulk MnSi antiferromagnetic [32, 33, 34, 35]. However, when grown as epitaxial thin films [16] the substrate stabilises the hexagonal unit cell of MnSi over the entire temperature range [15], fulfilling the symmetry requirements for making epitaxial MnSi a potential candidate for the altermagnetic phase with a d-wave symmetry. In good accordance with the theoretical predictions and the structural properties epitaxial MnSi exhibits a sizable spontaneous AHE despite a vanishingly small net magnetization [15].
In this work we investigate the anisotropy of AHE in epitaxial MnSi in terms of the external field orientation relative to the crystal structure. We show that a systematic rotation of the field in different crystal planes inflicts systematic changes on the magnetotransport properties, indicating that we are able to control the relative -orientation with experimentally reasonable external field strengths of the order of 2 T. Moreover, we show that changing the relative -orientation yields anisotropic behavior of the AHE, beyond cos, that requires a high crystal quality of the epitaxial films and is unlikely related to the magnetocrystalline anisotropy, which further demonstrates that the properties of epitaxial MnSi closely match those of altermagnetic materials.
II Methods
The epitaxial thin films of MnSi are grown on Si(111) substrates by molecular beam epitaxy using optimized growth parameters for realizing films with high crystalline quality and maximum MnSi content [16]. A few nm-thick MnSi phase acts as a thin seed layer aiding the nucleation of MnSi on Si(111) [16]. Unless otherwise specified, the films used in this work contain 96 % MnSi and 4 % MnSi and they have a thickness of 17 nm. The high crystallinity and the crystal orientation of the films have been verified using transmission electron microscopy (TEM) in our previous works [15, 16]. The epitaxial relationship between the Si substrate, MnSi seed layer and MnSi film is Si(111)//MnSi(111)//MnSi(0001). The films were patterned into microscopic Hall bars shown in Figure 1(b) using optical UV lithography and subsequent ion beam etching. The orientations of the Hall bar current channels relative to the crystal axis are defined by the angle as illustrated in Figure 1(d). We measure simultaneously the longitudinal () and transversal () voltages (denoted in Figure 1(b,c)) as a function of the external field strength and orientation to obtain the corresponding field- and angular-dependences of the longitudinal () and transversal () resistivities. We apply the external field always in the (yz) plane, i.e. the plane perpendicular to the current channel. Its exact orientation relative to the [0001] crystal axis (film normal) is given by the angle , as illustrated in Figure 1(c). Note that we will demonstrate later on in the text that the important parameter causing the anisotropic AHE is the crystal plane in which the external field rotates rather than the exact crystal orientation of the current channel - we alter the Hall bar orientation simply to ensure that the external field and the current channel are always perpendicular to each other unless otherwise stated. The current density is A cm throughout this work. For all the measurements presented in this work, the temperature was kept at 110 K, which is between the two critical temperatures of our films [15, 16]. At this temperature the films exhibit characteristic features of an altermagnetic phase, i.e. a sizable spontaneous AHE despite a vanishingly small net magnetization [15].
III Results and discussions
First, we study how the AHE resistivity () evolves when the external field orientation, parameterized by , is changed. We consider a Hall bar where the external field rotates in the (010) plane ( = 90 deg). We have first measured the hysteresis loops of for selected , from which a linear background (slope 0.045 cm ), originating from the ordinary Hall effect has been removed and only the antisymmetric-in-field part of the signal is used: . The symmetric part , is negligible in our sample at the specific measurement temperature supporting the collinear spin configuration [17]. The different contributions to are summarized in Figure 2(a). For the remaining of the paper, we will convert to AHE conductivity , which is used to characterize the intrinsic effects ( at 110 K is cm).
The hysteresis loops of for various are shown in Figure 2(b). We observe that the at saturation () is maximum when the field is close to the out-of-plane [0001] axis ( deg) and vanishes when the field is close to the in-plane [2] axis ( deg). As expected from the previous AHE study on the epitaxial MnSi [15], 180 degree rotation of this field changes the polarity of . The evolution of for the field orientations between [0001] and [000] is better visible in Figure 3 where we directly show as a function of for a 4 T external field. Between [0001] and [000] shows step-like behavior: when the field is oriented close to either [0001] or [2] (in-plane) crystal axes the slope is considerably more flat than in-between these directions. We note that the backward sweep does not trace the forward sweep because the saturation is not reached at 4 T at all field orientations as seen in Figure 2(b).
To unravel the possible origin of this anisotropic behavior of the AHE, we will first study the option of weak magnetization giving a contribution. Analogously to the AHE experiments, we have measured the sample magnetization along the external field for two field orientations and 90 deg as shown in Figures 2b and 3. After removing a linear diamagnetic background, we observe that the magnetization loops contain two components: a non-hysteretic part that relates to the substrate and the sample holder, and a hysteretic part that we attribute to the MnSi film as the magnitude of its coercivity agrees with that of the AHE-signal. This hysteretic component suggests a slight spontaneous canting of the magnetic sublattice moments and can play a role in the reversal mechanism of the Néel vector. We observe that the weak magnetization of the hysteretic component is small, i.e. u.c., and this value for and 90 deg is similar. We note that the behavior shown in Figure 2(b) would also be expected for a ferromagnet as at deg, while the wide plateau in Figure 3 could be a consequence of the magnetocrystalline anisotropy. We will now discuss further experiments to show that the anisotropic AHE is more likely to originate from the altermagnetic phase of our films.
First, we have measured the AHE in MnSi films with a larger proportion of MnSi, namely 75 % MnSi and 25 % MnSi. Increasing the MnSi content deteriorates the crystallinity of the MnSi [16], leading to a more textured film. We expect this to either prohibit altermagnetism as the strain necessary for stabilizing the hexagonal unit cell can relax at the grain boundaries, or even if the hexagonal unit cell is retained in the individual grains their different crystal orientations result in overall cancelling out of the altermagnetic effects. Figure 4 confirms that the magnetotransport properties of this film are different from those of the 96:4 MnSi:MnSi) MnSi epitaxial films with higher crystal quality. The field dependence of of the 75:25 control sample has a reduced amplitude and a vanishing coercivity, while the -dependence of the shows roughly a behavior, pointing to a ferromagnetic character. These results are in stark contrast with the behavior of the 96:4 sample and demonstrate that the AHE with anisotropic -dependence and non-zero coercivity requires a high crystal quality of the MnSi epitaxial films, which is consistent with the altermagnetic phase being at the origin of the AHE requiring the system to have a specific symmetry. Moreover, this shows that the spurious MnSi phase with 40 K is also unlikely origin of the AHE signal as the should in that case be stronger for the 75:25 sample.
Next, to further study the anisotropy of AHE, we vary systematically the crystal plane in which the external field and in effect the Néel vector (confirmed later using the anisotropic magnetoresistance measurements) is rotated and again observe how the is affected. This was realized by patterning multiple Hall bars with their current channels oriented along different crystal axes. We recall that these orientations are denoted by , which is the angle between the current channel and the crystal axis as was shown in Figure 1(d). For each Hall bar we rotate the field in a plane perpendicular to the current channel and thus effectively we are changing the crystal plane in which the field rotates. We have selected ’s that reflect the hexagonal crystal structure of our material and can be divided into two sets: for deg the current channel is along [] or equivalent so the field rotates in the crystal plane () or equivalent;
for deg the current channel is along [] or equivalent and the field rotates in the crystal plane () or equivalent. In Figure 5 we show the -dependence of the (with 4 T field) for the two sets of s. For each set, is maximum when the field is close to [0001] axis but significant differences arise when the field orients towards the sample plane: for deg, when the field orients in the [] direction or equivalent, we observe a wide plateau around = 90 deg, familiar from Figure 3, while for = 0, 60, 120 deg when the field orients in the [] direction or equivalent we observe a steeper linear slope instead. We have verified that this anisotropy persists also at 2 and 7 T fields. The hexagonal symmetry of the AHE signals for the different suggests that the three possibilities for the distribution of the magnetic Mn2 atoms (Figure 1) are likely present in equal populations: if each Mn2 site is on average magnetic, the -dependence of should reflect the hexagonal symmetry of the crystal structure. If one of these ’quadrupole domains’ dominated, lower symmetry for the -dependence of would be expected. We have verified that the key factor responsible for the change of behavior between the different s is the crystal plane in which the field rotates rather than the crystal axis along which the current channel is oriented, as observed in Figure 6. We note that for Figure 6 we have used a different sample of the same material, which explains some of the differences in the -dependence of the compared to the other figures. However, the overall behavior is the same, which confirms the reproducibility of the results discussed here. In summary, the results in Figures 5 and 6 allow us to conclude that the anisotropic behavior depends on the -orientation relative to the hexagonal crystal symmetry of the MnSi epitaxial films.
We also consider magnetocrystalline anisotropy as the origin of the anisotropic AHE unlikely: if we interpret the results shown in Figures 3 and 5 considering that the different crystal planes have strongly different magnetocrystalline energy landscapes, then for = 30, 90, and 150 deg it is more energetically favorable for to remain aligned along a crystal axis yielding at deg compared to = 0, 60, and 120 deg. However, the hysteresis loops measured at 90 deg for both 0 (not shown) and 90 deg (Figure 2(b)) superimpose at 4 T, indicating that it is unlikely that at this field is lagging behind the external field around 90 deg for = 30, 90, and 150 deg more than for = 0, 60, and 120 deg. Moreover, we observe the anisotropy in the -dependence of persisting up to 7 T, which is well above the saturation field of at any , as shown in Figure 2(b). Along with our demonstration that the anisotropic AHE is crystal quality-dependent and depends on the -orientation relative to the hexagonal crystal symmetry, these observations provide further support to the main AHE contribution in our MnSi films being from altermagnetism. These results constitute the main conclusion of our work and they systematically support the previous calculations and experiments on the same system [15].
Finally, we confirm the ability to manipulate with the external field by measuring the anisotropic magnetoresistance (AMR), which is known to depend on the -orientation. As shown in Figure 5, we have measured the - and -dependence of the longitudinal AMR (AMR), which relates to as AMR, where indicates averaged over all angles. AMR is an even function of the field and known to have two components: a non-crystalline one scaling with the angle between the order parameter and the current directions, and a crystalline one relating to the order parameter orientation relative to the crystal axes [36]. We show the -dependence for = 0, 60, 120 deg and = 30, 90, 150 deg Hall bars in Figure 5 at conditions identical to the AHE measurements. Again we observe different behaviors for the two sets of , as expected from the crystalline AMR depending on the crystal plane in which the order parameter rotates: for = 0, 60, 120 deg the dependence follows a behavior while for = 30, 90, 150 deg higher order even cosines appear in the signal.
Although the exact Néel vector orientation and rotation under the action of the external magnetic field cannot be deduced from the results presented, in this final paragraph we wish to open a discussion on the interpretation of the anisotropic AHE in the light of the symmetries of the epitaxial MnSi considering this material to be a strong candidate for the altermagnetic phase. First, it should be noted that in our measurement geometry (Figure 1), the changes in correspond to changes in the component of along the [0001] direction (out-of-plane). In altermagnets, the presence and orientation of depend on the -vector orientation because the latter influences the symmetries of the system. While accurate calculation of the entire -dependence of is inhibited by the unknown relevant micromagnetic parameters, the value of for some specific orientations can be inferred from symmetry-considerations. The symmetry analysis on the expected spin structure of the epitaxial MnSi shown in Figure 1(a) indicates that for non-zero the -vector cannot point along the [0001] axis ( excluded), along the main in-plane axes (only in-plane allowed), or in the planes of the high-symmetry in-plane axes highlighted in Figure 1(a) (only in-plane allowed). Here, the crystal symmetries of the planes corresponding to 0, 60, and 120 deg and 30, 90, and 150 deg differ, as seen in Figure 5 and therefore the likely differ, too.
IV Conclusion
In conclusion, the implications of our measurements on the external field orientation dependence of in epitaxial MnSi films are two-fold. First, the systematic changes in supported by the systematic changes in AMR upon the field rotation underlines our ability to manipulate the -orientation with an external field of the order of 2 T or more. This is an important premise for further experimentally feasible spin transport studies in altermagnets where control over the -orientation is critical. Second, our key result is that the in epitaxial MnSi films shows anisotropic behavior as a function of the external field-induced orientation of relative to the underlying crystal structure, and this anisotropic behavior depends on the crystal quality and is unlikely to relate to the magnetocrystalline anisotropy. This observation of the anisotropic adds to the growing array of properties of epitaxial MnSi advocating the altermagnetic phase in this material.
Acknowledgements.
This work was supported by the French national research agency (ANR) and the Deutsche Forschungsgemeinschaft (DFG) (Project MATHEEIAS- Grant No. ANR-20-CE92-0049-01 / DFG-445976410). H. R., E. S., S. T. B. G. and A. T. are supported by the DFG-GACR grant no. 490730630. J. R. acknowledges MINECO for the Margaritas Salas program. D. K. acknowledges the academy of sciences of the Czech Republic for the Lumina quaeruntur program.References
- Nagaosa et al. [2010] N. Nagaosa, J. Sinova, S. Onoda, A. H. MacDonald, and N. P. Ong, Anomalous Hall effect, Rev. Mod. Phys. 82, 1539 (2010).
- Kübler and Felser [2014] J. Kübler and C. Felser, Non-collinear antiferromagnets and the anomalous Hall effect, EPL (Europhysics Letters) 108, 67001 (2014).
- Chen et al. [2014] H. Chen, Q. Niu, and A. H. MacDonald, Anomalous Hall effect arising from noncollinear antiferromagnetism, Phys. Rev. Lett. 112, 017205 (2014).
- Nakatsuji et al. [2015] S. Nakatsuji, N. Kiyohara, and T. Higo, Large anomalous Hall effect in a non-collinear antiferromagnet at room temperature, Nature 527, 212 (2015).
- Nayak et al. [2016] A. K. Nayak, J. E. Fischer, Y. Sun, B. Yan, J. Karel, A. C. Komarek, C. Shekhar, N. Kumar, W. Schnelle, J. Kübler, C. Felser, and S. S. P. Parkin, Large anomalous Hall effect driven by a nonvanishing Berry curvature in the noncolinear antiferromagnet MGe, Science Advances 2, e1501870 (2016).
- Šmejkal et al. [2020] L. Šmejkal, R. González-Hernández, T. Jungwirth, and J. Sinova, Crystal time-reversal symmetry breaking and spontaneous Hall effect in collinear antiferromagnets, Science Advances 6, eaaz8809 (2020).
- Šmejkal et al. [2022] L. Šmejkal, A. H. MacDonald, J. Sinova, S. Nakatsuji, and T. Jungwirth, Anomalous Hall antiferromagnets, Nature Reviews Materials 7, 482 (2022).
- Jungwirth et al. [2016] T. Jungwirth, X. Marti, P. Wadley, and J. Wunderlich, Antiferromagnetic spintronics, Nature Nanotechnology 11, 231–241 (2016).
- Baltz et al. [2018] V. Baltz, A. Manchon, M. Tsoi, T. Moriyama, T. Ono, and Y. Tserkovnyak, Antiferromagnetic spintronics, Rev. Mod. Phys. 90, 015005 (2018).
- Šmejkal et al. [2022a] L. Šmejkal, J. Sinova, and T. Jungwirth, Beyond conventional ferromagnetism and antiferromagnetism: A phase with nonrelativistic spin and crystal rotation symmetry, Phys. Rev. X 12, 031042 (2022a).
- Feng et al. [2022] Z. Feng, X. Zhou, L. Šmejkal, L. Wu, Z. Zhu, H. Guo, R. González-Hernández, X. Wang, H. Yan, P. Qin, X. Zhang, H. Wu, H. Chen, Z. Meng, L. Liu, Z. Xia, J. Sinova, T. Jungwirth, and Z. Liu, An anomalous Hall effect in altermagnetic ruthenium dioxide, Nature Electronics 5, 735 (2022).
- Tschirner et al. [2023] T. Tschirner, P. Keßler, R. D. Gonzalez Betancourt, T. Kotte, D. Kriegner, B. Büchner, J. Dufouleur, M. Kamp, V. Jovic, L. Smejkal, J. Sinova, R. Claessen, T. Jungwirth, S. Moser, H. Reichlova, and L. Veyrat, Saturation of the anomalous hall effect at high magnetic fields in altermagnetic ruo2, APL Materials 11, 10.1063/5.0160335 (2023).
- Gonzalez Betancourt et al. [2023] R. D. Gonzalez Betancourt, J. Zubáč, R. Gonzalez-Hernandez, K. Geishendorf, Z. Šobáň, G. Springholz, K. Olejník, L. Šmejkal, J. Sinova, T. Jungwirth, S. T. B. Goennenwein, A. Thomas, H. Reichlová, J. Železný, and D. Kriegner, Spontaneous anomalous Hall effect arising from an unconventional compensated magnetic phase in a semiconductor, Phys. Rev. Lett. 130, 036702 (2023).
- Kluczyk et al. [2023] K. P. Kluczyk, K. Gas, M. J. Grzybowski, P. Skupiński, M. A. Borysiewicz, T. Fas, J. Suffczyński, J. Z. Domagala, K. Grasza, A. Mycielski, M. Baj, K. H. Ahn, K. Výborný, M. Sawicki, and M. Gryglas-Borysiewicz, Coexistence of anomalous Hall effect and weak net magnetization in collinear antiferromagnet MnTe (2023), arXiv:2310.09134 .
- Reichlová et al. [2021] H. Reichlová, R. L. Seeger, R. González-Hernández, I. Kounta, R. Schlitz, D. Kriegner, P. Ritzinger, M. Lammel, M. Leiviskä, V. Petříček, P. Doležal, E. Schmoranzerová, A. Bad’ura, A. Thomas, V. Baltz, L. Michez, J. Sinova, S. T. B. Goennenwein, T. Jungwirth, and L. Šmejkal, Macroscopic time reversal symmetry breaking by staggered spin-momentum interaction (2021), arXiv:2012.15651 [cond-mat.mes-hall] .
- Kounta et al. [2023] I. Kounta, H. Reichlova, D. Kriegner, R. Lopes Seeger, A. Bad’ura, M. Leiviska, A. Boussadi, V. Heresanu, S. Bertaina, M. Petit, E. Schmoranzerova, L. Smejkal, J. Sinova, T. Jungwirth, V. Baltz, S. T. B. Goennenwein, and L. Michez, Competitive actions of MnSi in the epitaxial growth of MnSi thin films on Si(111), Phys. Rev. Mater. 7, 024416 (2023).
- Badura et al. [2023] A. Badura, D. Kriegner, E. Schmoranzerová, K. Výborný, M. Leiviskä, R. L. Seeger, V. Baltz, D. Scheffler, S. Beckert, I. Kounta, L. Michez, L. Šmejkal, J. Sinova, S. T. B. Goennenwein, J. Železný, and H. Reichlová, Even-in-magnetic-field part of transverse resistivity as a probe of magnetic order (2023), arXiv:2311.14498 .
- Fedchenko et al. [2023] O. Fedchenko, J. Minar, A. Akashdeep, S. W. D’Souza, D. Vasilyev, O. Tkach, L. Odenbreit, Q. L. Nguyen, D. Kutnyakhov, N. Wind, L. Wenthaus, M. Scholz, K. Rossnagel, M. Hoesch, M. Aeschlimann, B. Stadtmueller, M. Klaeui, G. Schoenhense, G. Jakob, T. Jungwirth, L. Smejkal, J. Sinova, and H. J. Elmers, Observation of time-reversal symmetry breaking in the band structure of altermagnetic RuO (2023), arXiv:2306.02170 .
- Lee et al. [2023] S. Lee, S. Lee, S. Jung, J. Jung, D. Kim, Y. Lee, B. Seok, J. Kim, B. G. Park, L. Šmejkal, C.-J. Kang, and C. Kim, Broken Kramers’ degeneracy in altermagnetic MnTe (2023), arXiv:2308.11180 .
- Osumi et al. [2023] T. Osumi, S. Souma, T. Aoyama, K. Yamauchi, A. Honma, K. Nakayama, T. Takahashi, K. Ohgushi, and T. Sato, Observation of giant band splitting in altermagnetic MnTe (2023), arXiv:2308.10117 .
- Reimers et al. [2023] S. Reimers, L. Odenbreit, L. Smejkal, V. N. Strocov, P. Constantinou, A. B. Hellenes, R. J. Ubiergo, W. H. Campos, V. K. Bharadwaj, A. Chakraborty, T. Denneulin, W. Shi, R. E. Dunin-Borkowski, S. Das, M. Kläui, J. Sinova, and M. Jourdan, Direct observation of altermagnetic band splitting in CrSb thin films (2023), arXiv:2310.17280 .
- Sattigeri et al. [2023] R. M. Sattigeri, G. Cuono, and C. Autieri, Altermagnetic surface states: towards the observation and utilization of altermagnetism in thin films, interfaces and topological materials, Nanoscale 15, 16998–17005 (2023).
- Zhou et al. [2023] X. Zhou, W. Feng, R.-W. Zhang, L. Smejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal thermal transport in altermagnetic RuO (2023), arXiv:2305.01410 .
- González-Hernández et al. [2021] R. González-Hernández, L. Šmejkal, K. Výborný, Y. Yahagi, J. Sinova, T. c. v. Jungwirth, and J. Železný, Efficient electrical spin splitter based on nonrelativistic collinear antiferromagnetism, Phys. Rev. Lett. 126, 127701 (2021).
- Bose et al. [2022] A. Bose, N. J. Schreiber, R. Jain, D.-F. Shao, H. P. Nair, J. Sun, X. S. Zhang, D. A. Muller, E. Y. Tsymbal, D. G. Schlom, and D. C. Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nature Electronics 5, 267–274 (2022).
- Bai et al. [2022] H. Bai, L. Han, X. Y. Feng, Y. J. Zhou, R. X. Su, Q. Wang, L. Y. Liao, W. X. Zhu, X. Z. Chen, F. Pan, X. L. Fan, and C. Song, Observation of spin splitting torque in a collinear antiferromagnet RuO, Phys. Rev. Lett. 128, 197202 (2022).
- Chi et al. [2023] B. Chi, L. Jiang, Y. Zhu, G. Yu, C. Wan, J. Zhang, and X. Han, Crystal facet orientated altermagnets for detecting ferromagnetic and antiferromagnetic states by giant tunneling magnetoresistance effect (2023), arXiv:2309.09561 .
- Šmejkal et al. [2022b] L. Šmejkal, A. B. Hellenes, R. González-Hernández, J. Sinova, and T. Jungwirth, Giant and Tunneling magnetoresistance in unconventional collinear antiferromagnets with nonrelativistic spin-momentum coupling, Phys. Rev. X 12, 011028 (2022b).
- Xu et al. [2023] S. Xu, Z. Zhang, F. Mahfouzi, Y. Huang, H. Cheng, B. Dai, W. Cai, K. Shi, D. Zhu, Z. Guo, C. Cao, Y. Liu, A. Fert, N. Kioussis, K. L. Wang, Y. Zhang., and W. Zhao, Spin-flop magnetoresistance in a collinear antiferromagnetic tunnel junction (2023), arXiv:2311.02458 .
- Limmer et al. [2006] W. Limmer, M. Glunk, J. Daeubler, T. Hummel, W. Schoch, R. Sauer, C. Bihler, H. Huebl, M. S. Brandt, and S. T. B. Goennenwein, Angle-dependent magnetotransport in cubic and tetragonal ferromagnets: Application to (001)- and -oriented , Phys. Rev. B 74, 205205 (2006).
- Roman et al. [2009] E. Roman, Y. Mokrousov, and I. Souza, Orientation dependence of the intrinsic anomalous Hall effect in hcp cobalt, Phys. Rev. Lett. 103, 097203 (2009).
- Lander et al. [1967] G. H. Lander, P. J. Brown, and J. B. Forsyth, The antiferromagnetic structure of MnSi, Proceedings of the Physical Society 91, 332 (1967).
- Menshikov et al. [1990] A. Z. Menshikov, A. P. Vokhmyanin, and Y. A. Dorofeev, Magnetic structure and phase transformations in MnSi, physica status solidi (b) 158, 319 (1990).
- Brown et al. [1992] P. J. Brown, J. B. Forsyth, V. Nunez, and F. Tasset, The low-temperature antiferromagnetic structure of MnSi revised in the light of neutron polarimetry, Journal of Physics: Condensed Matter 4, 10025 (1992).
- Brown and Forsyth [1995] P. J. Brown and J. B. Forsyth, Antiferromagnetism in MnSi: the magnetic structure of the AF2 phase at 70 K, Journal of Physics: Condensed Matter 7, 7619 (1995).
- Rushforth et al. [2007] A. W. Rushforth, K. Výborný, C. S. King, K. W. Edmonds, R. P. Campion, C. T. Foxon, J. Wunderlich, A. C. Irvine, P. Vašek, V. Novák, K. Olejník, J. Sinova, T. Jungwirth, and B. L. Gallagher, Anisotropic magnetoresistance components in (Ga,Mn)As, Phys. Rev. Lett. 99, 147207 (2007).

### v2

@@
-Anisotropy of the anomalous Hall effect in the altermagnet candidate MnSi films
+Anisotropy of the anomalous Hall effect in the altermagnet candidate Mn5Si3 films
 Abstract
 Altermagnets are compensated magnets belonging to spin symmetry groups that allow alternating spin polarizations both in the coordinate space of the crystal and in the momentum space of the electronic structure.
-In these materials the anisotropic local crystal environment of the different sublattices lowers the symmetry of the system so that the opposite-spin sublattices are connected only by rotations, which results in an unconventional spin-polarized band structure in the momentum space.
+In these materials the anisotropic local crystal environment of the different sublattices lowers the symmetry of the system so that the opposite-spin sublattices are connected only by rotations.
+This results in an unconventional spin-polarized band structure in the momentum space.
 This low symmetry of the crystal structure is expected to be reflected in the anisotropy of the anomalous Hall effect.
-In this work, we study the anisotropy of the anomalous Hall effect in epitaxial thin films of MnSi, an altermagnetic candidate material.
+In this work, we study the anisotropy of the anomalous Hall effect in epitaxial thin films of Mn5Si3, an altermagnetic candidate material.
 We first demonstrate a change in the relative Néel vector orientation when rotating the external field orientation through systematic changes in both the anomalous Hall effect and the anisotropic longitudinal magnetoresistance.
-We then show that the anomalous Hall effect in this material is anisotropic with the Néel vector orientation relative to the crystal structure and that this anisotropy requires high crystal quality and unlikely correlates with the magnetocrystalline anisotropy.
-Our results provide further systematic support to the case for considering epitaxial thin films of MnSi as an altermagnetic candidate material.
+We then show that the anomalous Hall effect in this material is anisotropic with the Néel vector orientation relative to the crystal structure and that this anisotropy requires high crystal quality and unlikely stems from the magnetocrystalline anisotropy.
+Our results thus provide further systematic support to the case for considering epitaxial thin films of Mn5Si3 as an altermagnetic candidate material.
 I Introduction
@@
 In ferromagnetic materials the net magnetization is also a -odd axial vector and thus is allowed [1] while in magnetic materials with compensated order, is allowed upon sufficient lowering of the crystal symmetries [2, 3, 4, 5, 6, 7].
-In antiferromagnetic materials [8, 9] where the opposite-spin sublattices are connected by translation or inversion, is forbidden by symmetry, while in altermagnetic materials [10] this connecting symmetry is rotation, which allows for [11, 12, 13, 14, 15, 16, 17].
-This lowered symmetry in altermagnets stems from the two sublattices being inequivalent in terms of the local crystal environment.
-Altermagnets are characterised by alternating spin-splitting of the energy bands [18, 19, 20, 21, 22] and they show various core spintronic properties such as the AHE presented above, the Nernst effect (thermal counterpart of the AHE) [23], spin-polarized current generation [24, 25, 26], as well as giant and tunnel magnetoresistances [27, 28, 29] in heterostructures.
-In polycrystalline soft ferromagnets, is allowed by symmetry for any magnetization orientation.
-Therefore, the AHE is proportional to the component of the magnetization that is parallel to , i.e.
+In antiferromagnetic materials [8, 9] where the opposite-spin sublattices are connected by translation or inversion, is forbidden by symmetry, while in altermagnetic materials [10, 11, 12] this connecting symmetry is rotation, which can allow for although spin orientation along specific crystal axes can still forbid by symmetry [13, 14, 15, 16, 17, 18, 19, 20].
+We note that to observe AHE, a weak non-zero spin-orbit coupling is always required to transmit lattice information to the conduction electrons even though the origin of the high signal amplitude are the intrinsic symmetries and thus non-relativistic.
+The lowered symmetry in altermagnets stems from the two sublattices being inequivalent in terms of the local crystal environment.
+Altermagnets are characterised by alternating spin-splitting of the energy bands [10, 11, 21, 12, 22, 23, 24, 25, 26] and they show various core spintronic properties such as the AHE presented above, the Nernst effect (thermal counterpart of the AHE) [27, 28, 29], spin-polarized current generation [30, 31, 32, 33, 34], as well as giant and tunnel magnetoresistances [35, 36, 37] in heterostructures.
+In polycrystalline soft ferromagnets, is allowed by symmetry for any magnetization orientation and therefore the AHE is proportional to the component of the magnetization that is parallel to , i.e.
 to cos, where is the saturation magnetization and is the angle between the external field and the film normal, assuming that the field is strong enough to saturate the magnetization.
@@
 In altermagnets, specific Néel vector () orientations can either allow or exclude the Hall vector by symmetry [7, 13, 15] so that anisotropic AHE beyond cos can arise.
-Anisotropic AHE beyond cos in compensated magnets constitutes a hallmark of altermagnetism, as has been recently demonstrated in RuO [11] (intrinsic d-wave altermagnet, i.e.
+Anisotropic AHE beyond cos in compensated magnets constitutes a hallmark of altermagnetism, as has been recently demonstrated in RuO2 [13] (intrinsic d-wave altermagnet, i.e.
 the electronic structure displays four lobes with alternating spin-polarization) and MnTe showing spontaneous AHE [15] (intrinsic g-wave, which acquires a d-wave character in the presence of spin-orbit coupling).
-Both RuO and MnTe owe their altermagnetism to the presence of non-magnetic atoms that lower the symmetry of the system such that the opposite-spin sublattices are not connected by either inversion or translation [10].
-The saturation of in RuO and MnTe requires high fields of 68 [11, 12] and 4 T [13], respectively, so for further studies on the anisotropy of AHE in altermagnets a new material candidate where -saturation is realized at smaller fields would be ideal.
-Recently, epitaxial thin films of MnSi have emerged as such candidate [15] as the collinear compensated phase of MnSi with a hexagonal unit cell and four magnetic Mn atoms on the Mn2 sites in a checkerboard arrangement of the sublattices shown in Figure 1(a) fulfills the symmetry criteria for altermagnetism [15].
-Note, that in the bulk phase of MnSi the hexagonal unit cell is not preserved during the magnetic ordering and the opposite-spin sublattices become connected by translation making bulk MnSi antiferromagnetic [32, 33, 34, 35].
-However, when grown as epitaxial thin films [16] the substrate stabilises the hexagonal unit cell of MnSi over the entire temperature range [15], fulfilling the symmetry requirements for making epitaxial MnSi a potential candidate for the altermagnetic phase with a d-wave symmetry.
-In good accordance with the theoretical predictions and the structural properties epitaxial MnSi exhibits a sizable spontaneous AHE despite a vanishingly small net magnetization [15].
-In this work we investigate the anisotropy of AHE in epitaxial MnSi in terms of the external field orientation relative to the crystal structure.
+Both RuO2 and MnTe owe their altermagnetism to the presence of non-magnetic atoms that render the local crystal environments of the sublattices anisotropic [12].
+The saturation of in RuO2 and MnTe requires high fields of 68 [13, 14] and 4 T [15], respectively, so for further studies on the anisotropy of AHE in altermagnets a new material candidate where -saturation is realized at smaller fields would be ideal.
+Recently, high crystal quality thin films of Mn5Si3 have emerged as such candidate [17, 18, 19, 28, 29] as the collinear compensated phase of Mn5Si3 with a hexagonal unit cell and four magnetic Mn atoms on the Mn2 sites in a checkerboard arrangement of the sublattices (shown in Figure 1(a)) fulfills the symmetry criteria for altermagnetism [17].
+The altermagnetic phase has been supported by the experimental observations of both anomalous Hall [17, 19] and Nernst effects [28, 29].
+Note, that in the bulk phase of Mn5Si3 the hexagonal unit cell is not preserved during the magnetic ordering and the opposite-spin sublattices become connected by translation, making bulk Mn5Si3 antiferromagnetic [40, 41, 42, 43].
+However, when grown as epitaxial thin films [18] the substrate stabilizes the hexagonal unit cell of Mn5Si3 over the entire temperature range [17], fulfilling the symmetry requirements for making epitaxial Mn5Si3 a potential candidate for the altermagnetic phase with a d-wave symmetry.
+In good accordance with the theoretical predictions and the structural properties, epitaxial Mn5Si3 exhibits a sizable spontaneous AHE despite a vanishingly small net magnetization [17].
+We note that vanishing net magnetization and non-vanishing AHE are also observed in non-collinear antiferromagnets with broken time reversal symmetry, such as Mn3Sn and Mn3Ge [44].
+However, previous studies [17, 18, 19, 28, 29] through theoretical and experimental arguments show that non-collinear order in epitaxial Mn5Si3 films is unlikely.
+In this work we investigate the anisotropy of AHE in epitaxial Mn5Si3 in terms of the external field orientation relative to the crystal structure.
 We show that a systematic rotation of the field in different crystal planes inflicts systematic changes on the magnetotransport properties, indicating that we are able to control the relative -orientation with experimentally reasonable external field strengths of the order of 2 T.
-Moreover, we show that changing the relative -orientation yields anisotropic behavior of the AHE, beyond cos, that requires a high crystal quality of the epitaxial films and is unlikely related to the magnetocrystalline anisotropy, which further demonstrates that the properties of epitaxial MnSi closely match those of altermagnetic materials.
+Moreover, we show that changing the relative -orientation yields anisotropic behavior of the AHE, beyond cos, that requires a high crystal quality of the epitaxial films and is unlikely related to the magnetocrystalline anisotropy, which further demonstrates that the properties of epitaxial Mn5Si3 closely match those expected for altermagnetic materials.
+This is in line with previous findings in Mn5Si3 that predict collinear compensated hexagonal phase and demonstrate strong spontaneous anomalous Hall [17, 19] and Nernst [28, 29] effects with a clear influence of crystallinity [18] and Mn-content [29].
 II Methods
-The epitaxial thin films of MnSi are grown on Si(111) substrates by molecular beam epitaxy using optimized growth parameters for realizing films with high crystalline quality and maximum MnSi content [16].
-A few nm-thick MnSi phase acts as a thin seed layer aiding the nucleation of MnSi on Si(111) [16].
-Unless otherwise specified, the films used in this work contain 96 % MnSi and 4 % MnSi and they have a thickness of 17 nm.
+The epitaxial thin films of Mn5Si3 are grown on Si(111) substrates (thickness 250 m, 10 k cm) by molecular beam epitaxy using optimized growth parameters for realizing films with high crystalline quality and maximum Mn5Si3 content [18].
+A few nm-thick MnSi phase acts as a thin seed layer aiding the nucleation of Mn5Si3 on Si(111) [18].
+Unless otherwise specified, the films used in this work contain 96 % Mn5Si3 and 4 % MnSi and they have a thickness of 17 nm.
 The high crystallinity and the crystal orientation of the films have been verified using transmission electron microscopy (TEM) in our previous works [17, 18].
-The epitaxial relationship between the Si substrate, MnSi seed layer and MnSi film is Si(111)//MnSi(111)//MnSi(0001).
-The films were patterned into microscopic Hall bars shown in Figure 1(b) using optical UV lithography and subsequent ion beam etching.
-The orientations of the Hall bar current channels relative to the crystal axis are defined by the angle as illustrated in Figure 1(d).
-We measure simultaneously the longitudinal () and transversal () voltages (denoted in Figure 1(b,c)) as a function of the external field strength and orientation to obtain the corresponding field- and angular-dependences of the longitudinal () and transversal () resistivities.
+The epitaxial relationship between the Si substrate, MnSi seed layer and Mn5Si3 film is Si(111)//MnSi(111)//Mn5Si3(0001).
+The films were patterned into microscopic Hall bars shown in Figure 1(c) using optical UV lithography and subsequent ion beam etching.
+The orientations of the Hall bar current channels relative to the crystal axis are defined by the angle as illustrated in Figure 1(e).
+We measure simultaneously the longitudinal () and transversal () voltages (denoted in Figure 1(c,d)) as a function of the external field strength and orientation to obtain the corresponding field- and angular-dependences of the longitudinal () and transversal () resistivities.
 We apply the external field always in the (yz) plane, i.e.
 the plane perpendicular to the current channel.
-Its exact orientation relative to the [0001] crystal axis (film normal) is given by the angle , as illustrated in Figure 1(c).
+Its exact orientation relative to the [0001] crystal axis (film normal) is given by the angle , as illustrated in Figure 1(d).
 Note that we will demonstrate later on in the text that the important parameter causing the anisotropic AHE is the crystal plane in which the external field rotates rather than the exact crystal orientation of the current channel - we alter the Hall bar orientation simply to ensure that the external field and the current channel are always perpendicular to each other unless otherwise stated.
-The current density is A cm throughout this work.
+The current density is A cm-2 throughout this work.
 For all the measurements presented in this work, the temperature was kept at 110 K, which is between the two critical temperatures of our films [17, 18].
@@
 The different contributions to are summarized in Figure 2(a).
-For the remaining of the paper, we will convert to AHE conductivity , which is used to characterize the intrinsic effects ( at 110 K is cm).
+For the remaining of the paper, we will convert to AHE conductivity , which is used to characterize the intrinsic effects ( at 110 K is 154 cm).
 The hysteresis loops of for various are shown in Figure 2(b).
 We observe that the at saturation () is maximum when the field is close to the out-of-plane [0001] axis ( deg) and vanishes when the field is close to the in-plane [2] axis ( deg).
-As expected from the previous AHE study on the epitaxial MnSi [15], 180 degree rotation of this field changes the polarity of .
+As expected from the previous AHE study on the epitaxial Mn5Si3 [17], 180 degree rotation of this field changes the polarity of .
 The evolution of for the field orientations between [0001] and [000] is better visible in Figure 3 where we directly show as a function of for a 4 T external field.
@@
 To unravel the possible origin of this anisotropic behavior of the AHE, we will first study the option of weak magnetization giving a contribution.
-Analogously to the AHE experiments, we have measured the sample magnetization along the external field for two field orientations and 90 deg as shown in Figures 2b and 3.
-After removing a linear diamagnetic background, we observe that the magnetization loops contain two components: a non-hysteretic part that relates to the substrate and the sample holder, and a hysteretic part that we attribute to the MnSi film as the magnitude of its coercivity agrees with that of the AHE-signal.
+Magnetic characterization was carried out using a superconducting quantum interference device (SQUID) magnetometer (Quantum Design MPMS XL 5.0).
+Before measurement the sample was cleaned with acetone, alcohol, and deionized water, and them mounted (in two different orientations) on a sample holder made of two quartz capillaries with two dots of glue.
+Analogously to the AHE experiments, we have measured the sample magnetization along the external field for two field orientations and 90 deg as shown in Figures 2(b) and 3.
+After removing a linear diamagnetic background, we observe that the magnetization loops contain two components: a non-hysteretic part that relates to the substrate and the sample holder, and a hysteretic part that we attribute to the Mn5Si3 film as the magnitude of its coercivity agrees with that of the AHE-signal.
 This hysteretic component suggests a slight spontaneous canting of the magnetic sublattice moments and can play a role in the reversal mechanism of the Néel vector.
-We observe that the weak magnetization of the hysteretic component is small, i.e.
-u.c., and this value for and 90 deg is similar.
-We note that the behavior shown in Figure 2(b) would also be expected for a ferromagnet as at deg, while the wide plateau in Figure 3 could be a consequence of the magnetocrystalline anisotropy.
-We will now discuss further experiments to show that the anisotropic AHE is more likely to originate from the altermagnetic phase of our films.
-First, we have measured the AHE in MnSi films with a larger proportion of MnSi, namely 75 % MnSi and 25 % MnSi.
-Increasing the MnSi content deteriorates the crystallinity of the MnSi [16], leading to a more textured film.
-We expect this to either prohibit altermagnetism as the strain necessary for stabilizing the hexagonal unit cell can relax at the grain boundaries, or even if the hexagonal unit cell is retained in the individual grains their different crystal orientations result in overall cancelling out of the altermagnetic effects.
-Figure 4 confirms that the magnetotransport properties of this film are different from those of the 96:4 MnSi:MnSi) MnSi epitaxial films with higher crystal quality.
-The field dependence of of the 75:25 control sample has a reduced amplitude and a vanishing coercivity, while the -dependence of the shows roughly a behavior, pointing to a ferromagnetic character.
-These results are in stark contrast with the behavior of the 96:4 sample and demonstrate that the AHE with anisotropic -dependence and non-zero coercivity requires a high crystal quality of the MnSi epitaxial films, which is consistent with the altermagnetic phase being at the origin of the AHE requiring the system to have a specific symmetry.
+We convert the moment of the hysteretic component to magnetization using where the volume of the unit cell is , where a = 6.93 Å and c = 4.77 Å for our films at 110 K [17].
+The volume of the magnetic film is nm mm mm.
+We observe that the magnetization of the hysteretic component is u.c., and this value for and 90 deg is similar.
+While still being vanishingly small this value slightly exceeds the previously reported values [17, 19], which likely stems from the slight variations in the magnetic properties between the different thin films.
+We note that the non-zero AHE despite the vanishing net magnetization could also arise from non-collinear AFM ordering in our films [4, 5] but as previously shown this arrangement is unlikely in our material at the considered temperature [17].
+Alternatively, the behavior shown in Figure 2(b) would also be expected for a ferromagnet as at deg, while the wide plateau in Figure 3 could be a consequence of the magnetocrystalline anisotropy.
+We will now discuss further experiments to show that the anisotropic AHE is more likely to originate from the presumed [17, 18, 19, 28, 29] altermagnetic phase of our films.
+First, we have measured the AHE in Mn5Si3 films with a larger proportion of MnSi (Figure 4), namely 75 % Mn5Si3 and 25 % MnSi, but with a similar thickness of 16 nm and longitudinal resistivity of 190 cm at 110 K (for bare MnSi film, is 220 cm at 110 K).
+Increasing the MnSi content has been shown to deteriorate the crystallinity of the Mn5Si3 film, leading to a more textured film [18].
+In the extreme polycrystalline case, altermagnetism can be prohibited as the strain necessary for stabilizing the hexagonal unit cell can relax at the grain boundaries, or if the hexagonal unit cell is retained in the individual grains their different crystal orientations would result in overall cancelling out of the altermagnetic effects.
+In Figure 4(c) we show that the 75:25 sample is more mosaic than the 96:4 judging from its wider broadening of the full width at half maximum (FWHM) of the rocking curves around the Mn5Si3-(0002) reflections (shown in the inset of Figure 4).
+This increase in mosaicity seems to be sufficient to result in the magnetotransport properties of this film to be different from those of the 96:4 (Mn5Si3:MnSi) films with higher crystal quality: the field dependence of of the 75:25 control sample has a reduced amplitude and a vanishing coercivity, while the -dependence of the shows roughly a behavior, pointing to a ferromagnetic character.
+These results are in stark contrast with the behavior of the 96:4 sample and demonstrate that the AHE with anisotropic -dependence and non-zero coercivity requires a high crystal quality of the Mn5Si3 epitaxial films, which is consistent with the presumed [17, 18, 19, 28, 29] altermagnetic phase being at the origin of the AHE as it requires the system to have a specific symmetry.
 Moreover, this shows that the spurious MnSi phase with 40 K is also unlikely origin of the AHE signal as the should in that case be stronger for the 75:25 sample.
@@
 This was realized by patterning multiple Hall bars with their current channels oriented along different crystal axes.
-We recall that these orientations are denoted by , which is the angle between the current channel and the crystal axis as was shown in Figure 1(d).
-For each Hall bar we rotate the field in a plane perpendicular to the current channel and thus effectively we are changing the crystal plane in which the field rotates.
-We have selected ’s that reflect the hexagonal crystal structure of our material and can be divided into two sets: for deg the current channel is along [] or equivalent so the field rotates in the crystal plane () or equivalent;
+We recall that these orientations are denoted by , which is the angle between the current channel and the crystal axis as was shown in Figure 1(e).
+For each Hall bar we rotate the field in a plane perpendicular to the current channel and thus effectively change the crystal plane in which the field rotates.
+We have chosen to focus on only the high symmetry crystal axes selecting ’s that reflect the hexagonal crystal structure of our material and can be divided into two sets: for deg the current channel is along [] or equivalent so the field rotates in the crystal plane () or equivalent;
 for deg the current channel is along [] or equivalent and the field rotates in the crystal plane () or equivalent.
@@
 For each set, is maximum when the field is close to [0001] axis but significant differences arise when the field orients towards the sample plane: for deg, when the field orients in the [] direction or equivalent, we observe a wide plateau around = 90 deg, familiar from Figure 3, while for = 0, 60, 120 deg when the field orients in the [] direction or equivalent we observe a steeper linear slope instead.
-We have verified that this anisotropy persists also at 2 and 7 T fields.
-The hexagonal symmetry of the AHE signals for the different suggests that the three possibilities for the distribution of the magnetic Mn2 atoms (Figure 1) are likely present in equal populations: if each Mn2 site is on average magnetic, the -dependence of should reflect the hexagonal symmetry of the crystal structure.
-If one of these ’quadrupole domains’ dominated, lower symmetry for the -dependence of would be expected.
+Intermediate angles, like deg show intermediate behavior and provide no further information.
+We have verified that the anisotropy persists also at 2 and 7 T fields (Figure 7) .
+Moreover, the hexagonal symmetry of the AHE signals for the different suggests that the three possibilities for the quadrupole domains, which Néel vector may have a parallel component (Figure 1) are likely present in equal populations: if each Mn2 site is on average magnetic, the -dependence of should reflect the hexagonal symmetry of the crystal structure.
+If one of these quadrupole domains dominated, lower symmetry for the -dependence of would be expected.
+As the net AHE signal does not cancel out implies that the six magnetic domains corresponding to the three quadrupole domains and their reversed Néel vector counterparts are not equally populated.
+We have attempted to modify the population of the quadrupole domains through field-cooling in order to observe a corresponding break in the hexagonal symmetry of AHE signals.
+We field cooled the sample from 300 to 110 K along two different in-plane directions: [] and [] at 1.9 T field, after which we measured the angular -dependence at 110 K with a 1.9 T field, for and deg.
+We observed the same results as the zero-field cooled case considered in the manuscript (Figures 5,6, and 7), suggesting that the field cooling has no influence on the angular dependence of the AHE or the hexagonal anisotropy between the different crystal planes in which the field rotates.
+From this, we infer that the quadrupole domains could not be manipulated with field cooling.
+In fact, while the magnetic field will influence the Neel order, the response of the non-ordered Mn atoms (Figure 1) to the field is likely non-trivial and thus altering the population of the different quadrupole domains is also non-trivial.
+We also note that the presumed quadrupole domains in Mn5Si3 due to some Mn atoms not being ordered (Figure 1b) differ from other multipole domains, such as the cluster octupole domains in the Mn3Ge and Mn3Sn non-collinear antiferromagnets [44, 45], where all Mn atoms are magnetic and which can be manipulated with an external field.
 We have verified that the key factor responsible for the change of behavior between the different s is the crystal plane in which the field rotates rather than the crystal axis along which the current channel is oriented, as observed in Figure 6.
@@
 However, the overall behavior is the same, which confirms the reproducibility of the results discussed here.
-In summary, the results in Figures 5 and 6 allow us to conclude that the anisotropic behavior depends on the -orientation relative to the hexagonal crystal symmetry of the MnSi epitaxial films.
+In summary, the results in Figures 5 and 6 allow us to conclude that the anisotropic behavior of the AHE depends on the -orientation relative to the hexagonal crystal symmetry of the Mn5Si3 epitaxial films.
 We also consider magnetocrystalline anisotropy as the origin of the anisotropic AHE unlikely: if we interpret the results shown in Figures 3 and 5 considering that the different crystal planes have strongly different magnetocrystalline energy landscapes, then for = 30, 90, and 150 deg it is more energetically favorable for to remain aligned along a crystal axis yielding at deg compared to = 0, 60, and 120 deg.
-However, the hysteresis loops measured at 90 deg for both 0 (not shown) and 90 deg (Figure 2(b)) superimpose at 4 T, indicating that it is unlikely that at this field is lagging behind the external field around 90 deg for = 30, 90, and 150 deg more than for = 0, 60, and 120 deg.
-Moreover, we observe the anisotropy in the -dependence of persisting up to 7 T, which is well above the saturation field of at any , as shown in Figure 2(b).
-Along with our demonstration that the anisotropic AHE is crystal quality-dependent and depends on the -orientation relative to the hexagonal crystal symmetry, these observations provide further support to the main AHE contribution in our MnSi films being from altermagnetism.
+However, the hysteresis loops measured at 90 deg for both 0 and 90 deg (Figure 2(b)) superimpose at 4 T, indicating that it is unlikely that at this field is lagging behind the external field around 90 deg for = 30, 90, and 150 deg more than for = 0, 60, and 120 deg.
+Moreover, as shown in Figure 7 we observe the anisotropy in the -dependence of persisting up to 7 T, which is well above the saturation field of at any , as shown in Figure 2(b).
+Along with our demonstration that the anisotropic AHE is crystal quality-dependent and depends on the -orientation relative to the hexagonal crystal symmetry, these observations provide further support to the main AHE contribution in our Mn5Si3 films being from altermagnetism.
 These results constitute the main conclusion of our work and they systematically support the previous calculations and experiments on the same system [17].
 Finally, we confirm the ability to manipulate with the external field by measuring the anisotropic magnetoresistance (AMR), which is known to depend on the -orientation.
-As shown in Figure 5, we have measured the - and -dependence of the longitudinal AMR (AMR), which relates to as AMR, where indicates averaged over all angles.
-AMR is an even function of the field and known to have two components: a non-crystalline one scaling with the angle between the order parameter and the current directions, and a crystalline one relating to the order parameter orientation relative to the crystal axes [36].
+As shown in Figure 5, we have measured the - and -dependence of the longitudinal AMR (AMR||), which relates to as AMR, where indicates averaged over all angles.
+AMR is an even function of the field and known to have two components: a non-crystalline one scaling with the angle between the order parameter and the current direction, and a crystalline one relating to the order parameter orientation relative to the crystal axes [46].
 We show the -dependence for = 0, 60, 120 deg and = 30, 90, 150 deg Hall bars in Figure 5 at conditions identical to the AHE measurements.
 Again we observe different behaviors for the two sets of , as expected from the crystalline AMR depending on the crystal plane in which the order parameter rotates: for = 0, 60, 120 deg the dependence follows a behavior while for = 30, 90, 150 deg higher order even cosines appear in the signal.
-Although the exact Néel vector orientation and rotation under the action of the external magnetic field cannot be deduced from the results presented, in this final paragraph we wish to open a discussion on the interpretation of the anisotropic AHE in the light of the symmetries of the epitaxial MnSi considering this material to be a strong candidate for the altermagnetic phase.
+Although the exact Néel vector orientation and rotation under the action of the external magnetic field cannot be deduced from the results presented, we wish to open a discussion on the interpretation of the anisotropic AHE in the light of the symmetries of the epitaxial Mn5Si3.
+This material is considered to be a candidate for the altermagnetic phase [17, 18, 19, 28, 29].
 First, it should be noted that in our measurement geometry (Figure 1), the changes in correspond to changes in the component of along the [0001] direction (out-of-plane).
 In altermagnets, the presence and orientation of depend on the -vector orientation because the latter influences the symmetries of the system.
-While accurate calculation of the entire -dependence of is inhibited by the unknown relevant micromagnetic parameters, the value of for some specific orientations can be inferred from symmetry-considerations.
-The symmetry analysis on the expected spin structure of the epitaxial MnSi shown in Figure 1(a) indicates that for non-zero the -vector cannot point along the [0001] axis ( excluded), along the main in-plane axes (only in-plane allowed), or in the planes of the high-symmetry in-plane axes highlighted in Figure 1(a) (only in-plane allowed).
-Here, the crystal symmetries of the planes corresponding to 0, 60, and 120 deg and 30, 90, and 150 deg differ, as seen in Figure 5 and therefore the likely differ, too.
+While accurate calculation of the entire -dependence of is inhibited by the unknown relevant micromagnetic parameters, the value of for some specific orientations can be inferred from symmetry considerations.
+More specifically, considering the expected spin structure of the epitaxial Mn5Si3 shown in Figure 1(a) we can use symmetry analysis to obtain information on the -vector orientations that give zero in the xy direction: if -vector is along [0001] crystal axis, is forbidden in any direction due to , , and symmetries, while if -vector is along any high-symmetry in-plane axis (highlighted in Figure 1(a)), only in-plane is allowed due to the and symmetries.
+This illustrates the influence of the orientation on the through symmetry, and thus the fact that the symmetries of the planes corresponding to 0, 60, and 120 deg and 30, 90, and 150 deg are different (Figure 5) can give a plausible explanation to the anisotropy of in these planes.
 IV Conclusion
-In conclusion, the implications of our measurements on the external field orientation dependence of in epitaxial MnSi films are two-fold.
-First, the systematic changes in supported by the systematic changes in AMR upon the field rotation underlines our ability to manipulate the -orientation with an external field of the order of 2 T or more.
+In conclusion, the implications of our measurements on the external field orientation dependence of in epitaxial Mn5Si3 films are two-fold.
+First, the systematic changes in supported by the systematic changes in AMR|| upon the field rotation underlines our ability to manipulate the -orientation with an external field of the order of 2 T and above.
 This is an important premise for further experimentally feasible spin transport studies in altermagnets where control over the -orientation is critical.
-Second, our key result is that the in epitaxial MnSi films shows anisotropic behavior as a function of the external field-induced orientation of relative to the underlying crystal structure, and this anisotropic behavior depends on the crystal quality and is unlikely to relate to the magnetocrystalline anisotropy.
-This observation of the anisotropic adds to the growing array of properties of epitaxial MnSi advocating the altermagnetic phase in this material.
+Second, our key result is that the in epitaxial Mn5Si3 films shows anisotropic behavior as a function of the external field-induced orientation of relative to the underlying crystal structure, and this anisotropic behavior depends on the crystal quality and is unlikely to relate to the magnetocrystalline anisotropy.
+This observation of the anisotropic adds to the growing array of properties of epitaxial Mn5Si3: theoretical prediction of the collinear compensated hexagonal phase required and experimental demonstrations of vanishing net magnetization, strong spontaneous anomalous Hall [17, 19] and Nernst [28, 29] effects , and their dependence on the crystallinity- [18] and Mn-content[29].
+These properties together advocate the altermagnetic phase in this material.
 Acknowledgements.
@@
 K.
-acknowledges the academy of sciences of the Czech Republic for the Lumina quaeruntur program.References
+acknowledges the academy of sciences of the Czech Republic for the Lumina quaeruntur program.
+AB,ES,DK, LS,TJ acknowledge the Ministry of Education, Youth and Sports of the Czech Republic through the OP JAK call Excellent Research (TERAFIT Project No.
+CZ.02.01.01/00/22_008/0004594)References
 - Nagaosa et al. [2010] N. Nagaosa, J. Sinova, S. Onoda, A. H. MacDonald, and N. P. Ong, Anomalous Hall effect, Rev. Mod. Phys. 82, 1539 (2010).
@@
 P.
-Parkin, Large anomalous Hall effect driven by a nonvanishing Berry curvature in the noncolinear antiferromagnet MGe, Science Advances 2, e1501870 (2016).
+Parkin, Large anomalous Hall effect driven by a nonvanishing Berry curvature in the noncolinear antiferromagnet M3Ge, Science Advances 2, e1501870 (2016).
 - Šmejkal et al.
@@
 - Baltz et al. [2018] V. Baltz, A. Manchon, M. Tsoi, T. Moriyama, T. Ono, and Y. Tserkovnyak, Antiferromagnetic spintronics, Rev. Mod. Phys. 90, 015005 (2018).
+- Hayami et al. [2019] S. Hayami, Y. Yanagi, and H. Kusunose, Momentum-dependent spin splitting by collinear antiferromagnetic ordering, Journal of the Physical Society of Japan 88, 123702 (2019).
+- Yuan et al. [2020] L.-D. Yuan, Z. Wang, J.-W. Luo, E. I. Rashba, and A. Zunger, Giant momentum-dependent spin splitting in centrosymmetric low- antiferromagnets, Phys. Rev. B 102, 014422 (2020).
 - Šmejkal et al.
@@
 Reichlova, and L.
-Veyrat, Saturation of the anomalous hall effect at high magnetic fields in altermagnetic ruo2, APL Materials 11, 10.1063/5.0160335 (2023).
+Veyrat, Saturation of the anomalous hall effect at high magnetic fields in altermagnetic ruo2, APL Materials 11, 101103 (2023).
 - Gonzalez Betancourt et al.
@@
 Jungwirth, and L.
-Šmejkal, Macroscopic time reversal symmetry breaking by staggered spin-momentum interaction (2021), arXiv:2012.15651 [cond-mat.mes-hall] .
+Šmejkal, Macroscopic time reversal symmetry breaking by staggered spin-momentum interaction (2021), arXiv:2012.15651 .
 - Kounta et al.
@@
 Goennenwein, and L.
-Michez, Competitive actions of MnSi in the epitaxial growth of MnSi thin films on Si(111), Phys.
+Michez, Competitive actions of MnSi in the epitaxial growth of Mn5Si3 thin films on Si(111), Phys.
 Rev.
@@
 7, 024416 (2023).
+- Han et al.
+[2024a] L.
+Han, X.
+Fu, R.
+Peng, X.
+Cheng, J.
+Dai, L.
+Liu, Y.
+Li, Y.
+Zhang, W.
+Zhu, H.
+Bai, Y.
+Zhou, S.
+Liang, C.
+Chen, Q.
+Wang, X.
+Chen, L.
+Yang, Y.
+Zhang, C.
+Song, J.
+Liu, and F.
+Pan, Electrical 180° switching of néel vector in spin-splitting antiferromagnet, Science Advances 10, eadn0479 (2024a).
 - Badura et al.
@@
 Reichlová, Even-in-magnetic-field part of transverse resistivity as a probe of magnetic order (2023), arXiv:2311.14498 .
+- Mazin et al.
+[2021] I.
+I.
+Mazin, K.
+Koepernik, M.
+D.
+Johannes, R.
+González-Hernández, and L.
+Šmejkal, Prediction of unconventional magnetism in doped fesb 2, Proceedings of the National Academy of Sciences 118, 10.1073/pnas.2108924118 (2021).
 - Fedchenko et al.
@@
 J.
-Elmers, Observation of time-reversal symmetry breaking in the band structure of altermagnetic RuO (2023), arXiv:2306.02170 .
+Elmers, Observation of time-reversal symmetry breaking in the band structure of altermagnetic RuO2, Science Advances 10, eadj4883 (2024).
 - Lee et al.
@@
 Kang, and C.
-Kim, Broken Kramers’ degeneracy in altermagnetic MnTe (2023), arXiv:2308.11180 .
+Kim, Broken kramers degeneracy in altermagnetic mnte, Phys.
+Rev.
+Lett.
+132, 036702 (2024).
 - Osumi et al.
@@
 Ohgushi, and T.
-Sato, Observation of giant band splitting in altermagnetic MnTe (2023), arXiv:2308.10117 .
+Sato, Observation of a giant band splitting in altermagnetic mnte, Phys.
+Rev.
+B 109, 115102 (2024).
 - Reimers et al.
@@
 Sinova, and M.
-Jourdan, Direct observation of altermagnetic band splitting in CrSb thin films (2023), arXiv:2310.17280 .
+Jourdan, Direct observation of altermagnetic band splitting in CrSb thin films, Nature Communications 15, 2116 (2024).
 - Sattigeri et al.
@@
 Cuono, and C.
-Autieri, Altermagnetic surface states: towards the observation and utilization of altermagnetism in thin films, interfaces and topological materials, Nanoscale 15, 16998–17005 (2023).
-- Zhou et al. [2023] X. Zhou, W. Feng, R.-W. Zhang, L. Smejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal thermal transport in altermagnetic RuO (2023), arXiv:2305.01410 .
-- González-Hernández et al.
-[2021] R.
-González-Hernández, L.
-Šmejkal, K.
-Výborný, Y.
-Yahagi, J.
+Autieri, Altermagnetic surface states: towards the observation and utilization of altermagnetism in thin films, interfaces and topological materials, Nanoscale 15, 16998 (2023).
+- Zhou et al. [2024] X. Zhou, W. Feng, R.-W. Zhang, L. Šmejkal, J. Sinova, Y. Mokrousov, and Y. Yao, Crystal thermal transport in altermagnetic , Phys. Rev. Lett. 132, 056701 (2024).
+- Badura et al.
+[2024] A.
+Badura, W.
+H.
+Campos, V.
+K.
+Bharadwaj, I.
+Kounta, L.
+Michez, M.
+Petit, J.
+Rial, M.
+Leiviskä, V.
+Baltz, F.
+Krizek, D.
+Kriegner, J.
+Zemen, S.
+Telkamp, S.
+Sailler, M.
+Lammel, R.
+J.
+Ubiergo, A.
+B.
+Hellenes, R.
+González-Hernández, J.
 Sinova, T.
-c.
-v.
-Jungwirth, and J.
-Železný, Efficient electrical spin splitter based on nonrelativistic collinear antiferromagnetism, Phys.
-Rev.
-Lett.
-126, 127701 (2021).
+Jungwirth, S.
+T.
+B.
+Goennenwein, L.
+Šmejkal, and H.
+Reichlova, Observation of the anomalous nernst effect in altermagnetic candidate mn5si3 (2024), arXiv:2403.12929 .
+- Han et al.
+[2024b] L.
+Han, X.
+Fu, W.
+He, Y.
+Zhu, J.
+Dai, W.
+Yang, W.
+Zhu, H.
+Bai, C.
+Chen, C.
+Wan, X.
+Han, C.
+Song, J.
+Liu, and F.
+Pan, Observation of non-volatile anomalous nernst effect in altermagnet with collinear néel vector (2024b), arXiv:2403.13427 .
+- Naka et al. [2019] M. Naka, S. Hayami, H. Kusunose, Y. Yanagi, Y. Motome, and H. Seo, Spin current generation in organic antiferromagnets, Nature Communications 10, 4305 (2019).
+- Ma et al.
+[2021] H.-Y.
+Ma, M.
+Hu, N.
+Li, J.
+Liu, W.
+Yao, J.-F.
+Jia, and J.
+Liu, Multifunctional antiferromagnetic materials with giant piezomagnetism and noncollinear spin current, Nature Communications 12, 4305 (2021).
 - Bose et al.
@@
 C.
-Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nature Electronics 5, 267–274 (2022).
+Ralph, Tilted spin current generated by the collinear antiferromagnet ruthenium dioxide, Nature Electronics 5, 267 (2022).
 - Bai et al.
@@
 Fan, and C.
-Song, Observation of spin splitting torque in a collinear antiferromagnet RuO, Phys.
+Song, Observation of spin splitting torque in a collinear antiferromagnet RuO2, Phys.
 Rev.
@@
 128, 197202 (2022).
+- Karube et al.
+[2022] S.
+Karube, T.
+Tanaka, D.
+Sugawara, N.
+Kadoguchi, M.
+Kohda, and J.
+Nitta, Observation of spin-splitter torque in collinear antiferromagnetic , Phys.
+Rev.
+Lett.
+129, 137201 (2022).
 - Chi et al.
@@
 Zhang, and X.
-Han, Crystal facet orientated altermagnets for detecting ferromagnetic and antiferromagnetic states by giant tunneling magnetoresistance effect (2023), arXiv:2309.09561 .
+Han, Crystal-facet-oriented altermagnets for detecting ferromagnetic and antiferromagnetic states by giant tunneling magnetoresistance, Phys.
+Rev.
+Appl.
+21, 034038 (2024).
 - Šmejkal et al.
@@
 - Roman et al. [2009] E. Roman, Y. Mokrousov, and I. Souza, Orientation dependence of the intrinsic anomalous Hall effect in hcp cobalt, Phys. Rev. Lett. 103, 097203 (2009).
-- Lander et al. [1967] G. H. Lander, P. J. Brown, and J. B. Forsyth, The antiferromagnetic structure of MnSi, Proceedings of the Physical Society 91, 332 (1967).
-- Menshikov et al. [1990] A. Z. Menshikov, A. P. Vokhmyanin, and Y. A. Dorofeev, Magnetic structure and phase transformations in MnSi, physica status solidi (b) 158, 319 (1990).
+- Lander et al. [1967] G. H. Lander, P. J. Brown, and J. B. Forsyth, The antiferromagnetic structure of Mn5Si3, Proceedings of the Physical Society 91, 332 (1967).
+- Menshikov et al. [1990] A. Z. Menshikov, A. P. Vokhmyanin, and Y. A. Dorofeev, Magnetic structure and phase transformations in Mn5Si3, physica status solidi (b) 158, 319 (1990).
 - Brown et al.
@@
 Nunez, and F.
-Tasset, The low-temperature antiferromagnetic structure of MnSi revised in the light of neutron polarimetry, Journal of Physics: Condensed Matter 4, 10025 (1992).
-- Brown and Forsyth [1995] P. J. Brown and J. B. Forsyth, Antiferromagnetism in MnSi: the magnetic structure of the AF2 phase at 70 K, Journal of Physics: Condensed Matter 7, 7619 (1995).
+Tasset, The low-temperature antiferromagnetic structure of Mn5Si3 revised in the light of neutron polarimetry, Journal of Physics: Condensed Matter 4, 10025 (1992).
+- Brown and Forsyth [1995] P. J. Brown and J. B. Forsyth, Antiferromagnetism in Mn5Si3: the magnetic structure of the AF2 phase at 70 K, Journal of Physics: Condensed Matter 7, 7619 (1995).
+- Chen et al.
+[2021] T.
+Chen, T.
+Tomita, S.
+Minami, M.
+Fu, T.
+Koretsune, M.
+Kitatani, I.
+Muhammad, D.
+Nishio-Hamane, R.
+Ishii, F.
+Ishii, R.
+Arita, and S.
+Nakatsuji, Anomalous transport due to weyl fermions in the chiral antiferromagnets mn3x, x=sn, ge, Nature Communications 12, 572 (2021).
+- Sugimoto et al.
+[2020] S.
+Sugimoto, Y.
+Nakatani, Y.
+Yamane, M.
+Ikhlas, K.
+Kondou, M.
+Kimata, T.
+Tomita, S.
+Nakatsuji, and Y.
+Otani, Electrical nucleation, displacement, and detection of antiferromagnetic domain walls in the chiral antiferromagnet mn3sn, Communications Physics 3, 111 (2020).
 - Rushforth et al.